### v1

@@ 4.3 Challenges of LLMs in Generating Reliable and Accurate Answers Across Specific Domains
RQ3 answer: Although there are structural differences between ChatGPT and LLaMA-2, the LLMs partially share the same challenges. Questions about general conventions or unpopular features of Framework and Libraries pose a challenge for the LLMs evaluated here, while Programming Language questions are expected to be addressed appropriately. The number of tags associated with a question negatively correlates with textual similarity provided by ChatGPT. Although ChatGPT presents a more neutral sentiment when generating their answers, LLaMA stands out with a more positive behavior.
4.4 ChatGPT Impact on User Activeness and Arranging it with Stack Overflow to solve questions
@@
Active Users
Pre-Release 314,179 62,007 67,240 239,045
Active Users
@@
Inactive Users
Post-Release 259,981 49,457 56,671 179,300
Exclusively Active
Users Post-Release 239,613 43,967 46,876 146,785
Reminiscent Users
Post-Release 54,198 12,550 10,569 59,745
Table 2: User activeness on Stack Overflow. The second and third rows show the number of users before and after ChatGPT release, respectively. The fourth row presents the inactive users (the difference between the second and third rows), while the fifth row shows the active users exclusively after the release (the difference between the third and second rows). Finally, the sixth row presents the reminiscent users (intersection between the second and third rows).
@@
In this research question, we further investigate the impact of user engagement on Stack Overflow after ChatGPT release, but focusing on the user side. Table 2 presents the general observed frequency of users on Stack Overflow. In conformity with our previous findings, we observe a general drop in active users (2nd and 3rd rows, respectively, in Table 2). Regardless of the type of user (questioner, and/or respondent, and commentator), all present a consistent drop. Such observation corroborates the drop in content posted on Stack Overflow as well (see Section 4.1).
When we compare the users active after the ChatGPT release, based on the same previous categories, we observe that most users were not active before (4th row in Table 2). Some may argue that these new active users could be associated with new accounts created after the release. However, when checking the creation date of these accounts, most of them were created before ChatGPT release (see Table 3). Since different factors might influence this observation, like sporadic users and exploring new sources of assistance, like LLMs, it is important to highlight the commitment users have to their community. Although we report a drop in the number of posted questions on Stack Overflow, these active users posting their new questions contribute to new data being generated and, consequently, new versions of LLMs being trained on them. In Section 5, we further discuss the impact and the need for generating data and LLMs.
Considering the challenges shared by LLMs previously discussed in RQ3 (see Section 4.3), we further explore the different types of users, aiming to understand whether and how they face the previous challenges. For that, we manually analyze questions that were possibly asked to ChatGPT, while they keep a reference for such event in its contents (question title or body). Initially, by analyzing all the 578,115 questions from this study (see Table 1), which also include questions without answers, we observe that 2193 questions mention ChatGPT in their contents (title or body). Next, checking the creation date of the questioners’ accounts, we observe that 1439 questions were asked by reminiscent users, while 731 questions were asked by new ones. The remaining questions were asked in anonymous mode or by users no longer with active accounts in Stack Overflow, so we could not check whether the questions were associated with reminiscent or new users. From each reminiscent and new users group, we randomly select a subsample of questions to be manually analyzed (91 and 86 questions, respectively), aiming for a 95% confidence level and a 10% margin of error. Finally, we also adopt the same approach and method used for RQ3 (see Section 4.3), with further analysis as we explain as follows.
First, to investigate whether users replicate questions on Stack Overflow, which were previously asked on ChatGPT, we checked whether the users explicitly mentioned that statement in the question content. Overall, users adopt this behavior for more than 87% of all analyzed questions (82% and 93% for reminiscent and new users, respectively). The remaining questions ask for clarification about the usage of ChatGPT, like its API. Further exploring these questions, though users mention they have consulted ChatGPT for clarifications, 60% of the questions do not present the answers generated by ChatGPT (49% and 70% for reminiscent and new users, respectively). The main reason pointed out by the users is that the answers provided do not fix their problems, as they have tried different proposed suggestions. For example, when a user asks about constructing an imbalanced dataset based on a pre-defined gini-coefficient, they report that the ChatGPT could not even generate a working function. [^18] https://stackoverflow.com/questions/76008808/how-to-construct-an-imbalanced-mnist-dataset-based-on-a-pre-defined-gini-coeffic. However, for the users that consider the suggestions provided by ChatGPT useful, they inform them as a starting point for further discussions. For example, in one question, a user asks for support when using the Overpass API. [^19] https://stackoverflow.com/questions/75558305/get-adjacent-ways-overpass-api Based on the initial solution provided, the respondent analyzed it and provided their suggestion, which, based on the respondent, was better than the solution provided by ChatGPT.
@@
Regarding the topics that represent challenges for Stack Overflow users, we observe that reminiscent and new users face the same challenges when trying to get answers from ChatGPT, and later appeal to Stack Overflow. Similar to the challenge domains that LLMs face reported in RQ3 (see Section 4.3), we observe that questions regarding Frameworks and Libraries represent the most recurrent challenge faced by reminiscent and new users (41% and 38%, respectively), followed by question targeting Programming Languages (22% and 25%, for reminiscent and new users, respectively). Besides the hardness of addressing questions targeting these topics by LLMs, users are aware of these limitations as they recognize these issues. For example, when a user asks for support regarding starting a new project using multiple frameworks like Next.JS, they consider the provided support of ChatGPT was not very helpful. They further argue that the failed ChatGPT’s attempt is due to their limited access to information dated until late 2021.202020https://stackoverflow.com/questions/75861694/useful-boilerplate-for-nextjs-nextauth-cognito-and-hasura
@@
RQ4 answer: The introduction of ChatGPT also led to a drop in the number of active users in Stack Overflow, even under different roles (questioners, respondents, and commentators). Most active users’ pos-release were inactive before ChatGPT, showing that they are reminiscent users, as their accounts were created before the release. Finally, reminiscent and new users share the same challenges when addressing their questions on ChatGPT and later appealing to Stack Overflow.
4.5 Exploring the Individual Impact of ChatGPT on Different Domains
Considering the results of our previous RQs, we further explore here the impact of ChatGPT on different domains associated with questions on Stack Overflow. For that, based on the domains that might (i) represent a challenge for LLMs (RQ3) and (ii) those that users might appeal to Stack Overflow to get support for them (RQ4), we statistically compare the number of questions posted before and after the release. This way, we explore questions associated with Programming Languages and Frameworks and Libraries. Although we have observed a general drop in posted content in RQ1, we believe that for some topics, no impact might have affected them. For each domain analyzed here, we first select the ten most cited tags before the release of ChatGPT. Although Frameworks and Libraries are classified as one single domain, we explore them individually, selecting the ten most cited frameworks and libraries. First, we discuss our findings for programming languages, followed by frameworks and libraries.
Figure 7 shows the distribution of questions associated with the top 10 most cited programming languages. Overall, we can observe that Python is the most common programming language required by users, followed by JavaScript. Such results might be motivated by the continuous adoption of these languages for Machine Learning and Web Development, respectively. The remaining programming languages present similar frequencies, even oscillating positions with each other, but with a large difference from the top two languages (placed on the bottom of the chart).
Analyzing the chart, we can observe a clear decline in posted questions after the release of ChatGPT for Python and JavaScript (top 2), the same observation made for general posted content analyzed in RQ1. However, for the remaining programming languages, there is no clear drop as they present a similar frequency of Posted questions before and after the release. Statically assessing whether there is a significant difference between the number of posted questions through the Wilcoxon rank sum test, we can reject the null hypothesis for all the evaluated programming languages (p-value < 0.01).
Different Impact on Most Cited Frameworks and Libraries
Figures 8 and  9 show the distribution of questions associated with the top ten most cited frameworks and libraries on Stack Overflow, respectively. Different from the chart for programming languages (Figure 7), we can observe different patterns regarding the distribution of questions here. Regarding the analysis of frameworks, we observe that ReactJS is the most cited framework, largely isolated from the others. Furthermore, that is the only framework in which we can clearly see a drop regarding the distribution of questions. Next, we have NodeJS and Android placed in the second and third positions, respectively. Although they oscillate their positions, Android is consistently in the second position for most of the evaluated time. Finally, the remaining frameworks are placed at the bottom of the chart, with some oscillations over time. Overall, it is also important to highlight the massive presence of frameworks based on JavaScript.
A similar distribution is also observed for the analysis of libraries. As highlighted in Figure 9, pandas is the most cited library, placed at the top of the chart with a large distance from the others. As previously discussed for the programming language Python, the adoption of pandas might be related to their adoption in Machine Learning, as it is commonly used for different tasks, like analyzing, cleaning, exploring, and manipulating data. Next, we observe a group of four libraries placed in the middle of the chart, while the remaining ones are placed in the bottom, oscillating positions over time. Finally, similar to the observations made for the framework analysis, here we observe a consistent number of libraries supporting different tasks on Machine Learning topics, showing the activeness of users on these topics.
When statistically analyzing the difference between the number of posted questions through the Wilcoxon rank sum test, we observe different results when compared to our analysis of programming languages. Regarding the frameworks, we reject the null hypothesis for eight frameworks except for Spring Boot and Angular (p-value < 0.01)212121p-value = 0.06 and 0.04, respectively. A similar behavior was observed for the libraries under analysis, as we reject the null hypothesis for all of them except ggplot2 (p-value = 0.34). Although a small fraction of subjects did not present a statistical drop in posted questions, such results corroborate our previous findings related to unsupported/challenged domains of Stack Overflow on ChatGPT.
Different factors might contribute to this particular result. Considering the evolution of these frameworks and libraries, LLMs are currently limited by the data they were trained on. As previously discussed in RQ4 (see Section 4.4), LLMs can not support tasks that involve information generated after their training. This way, as discussed in RQ4, users still appeal to Stack Overflow to get support for their questions. In Section 5, we discuss the impact and the importance of still continuing to generate data for future LLMs’ versions. Another possible reason is related to the fact that different communities associated with these topics desire to keep active by supporting their members. Even in such cases, though data generation does not represent a challenge here, keeping these communities active and supportive is a challenge that might be faced.
RQ5 answer: Despite the observed general drop in posted content in Stack Overflow, the introduction of ChatGPT has different impacts on different topics (domains). Even analyzing the most cited topics in different domains, we observe that, in some cases, there is no statistical difference regarding the posted contents. This way, it is important to investigate the motivation and related challenges for keeping communities active and supportive on Stack Overflow.

### v2

@@ 1 Map <String, String> h;
RQ2 answer: Although there are structural differences between ChatGPT and LLaMA-2, the LLMs partially share the same challenges. Questions about general conventions or unpopular features of Framework and Libraries pose a challenge for the LLMs evaluated here, while Programming Language questions are expected to be addressed appropriately. Additional material linked to the questions might play a role in the similarity reached by the LLMs, like code snippets with additional related information. The number of tags associated with a question negatively correlates with textual similarity provided by ChatGPT.
4.3 RQ3: What is the impact of ChatGPT ’s public release on Stack Overflow posting activity?
(a) Questions
(b) Answers
(c) Comments
Figure 6: Violin plots comparing the distribution of the data across three time intervals: Before (July to November 2022), After (November 2023 to May 2023), and One-Year Milestone (November 2023 to May 2024).
(a) Questions
(b) Answers
(c) Comments
Figure 7: Distribution of posted content on Stack Overflow (questions, answers, and comments). The gap between the After and One-Year Milestone intervals represents the period from May to November 2023, during which no data was collected.
Our third research question investigates the impact of ChatGPT’s release on the content posted to Stack Overflow. First, we discuss the overall impact of posting and related activities, and then we discuss the impact on specific domains.
Impact on Posting Activities in Stack Overflow
To address this research question, we first collected the relevant data, further comparing their frequency regarding the number of posted questions, answers, and comments before and after the introduction of ChatGPT (see Table 1). Notably, the number of submitted answers exhibits the most significant decline: 543 533543533543\,533543 533 answers before and 425 391425391425\,391425 391 answers after ChatGPT (i.e., -22%). After the one-year milestone, the decline is even more significant, with a drop of 52% (from 425 391425391425\,391425 391 to 202 326202326202\,326202 326 answers). This decline has a subsequent impact on the number of questions with accepted answers, which dropped from 212 775212775212\,775212 775 before ChatGPT to 157 187157187157\,187157 187 after ChatGPT, and consequently to 78 2627826278\,26278 262 after one year of the release (i.e., -26% and -50%, respectively). Comments and questions follow closely, showing a slightly more significant decline after one year of the release (46% and 48%, respectively).
To gain initial insights into potential trends regarding the decline, Figures 6 and 7 illustrate the content and temporal evolution of Stack Overflow. Figure 6 reveals that before ChatGPT’s release, there was a higher density of posted content in the upper region of the violin plots. Just after the release, this concentration appears more evenly distributed across the plots, including the middle and lower regions. Finally, after the one-year milestone, we observe a drastic decrease in the concentration of contents. Overall, the median daily count of submitted content is lower after ChatGPT’s introduction. Figure 7 shows the decline in posted content over time. Importantly, oscillating patterns present in the data before ChatGPT’s release continue to be observed (they correspond to weekends).
[…]
RQ3 answer: After the introduction of ChatGPT, we observe a statistically significant decline in users’ activity on Stack Overflow. This is evidenced by the reduced frequency of posted questions, answers, and comments usage, with the most pronounced decline observed after the one-year milestone following ChatGPT’s release. Hence, ChatGPT’s introduction has impacted the platform’s dynamics, likely due to its ability to address many of the queries previously posted on Stack Overflow.
4.3.1 RQ3.1: How are different domains from Stack Overflow impacted by the release of ChatGPT?
To better understand the impact of ChatGPT’s release in different domains, based on the domains that represent a challenge for LLMs (RQ2), we aim to statistically compare the number of questions posted in each domain. This way, we explore questions associated with Programming Languages and Frameworks and Libraries. Figure 8 presents the trend analysis for the Top-1 tags of each domain evaluated here. Analyzing the charts, we observe that the number of mentions for tags is generally decreasing as time progresses, supporting our previous findings when we observed a general drop in posted content. However, we still aim to investigate each domain individually, as they might suffer different impacts. First, we discuss our findings for programming languages, followed by frameworks and libraries.
Figure 9 shows the distribution of questions associated with the top 10 most cited programming languages. Overall, we can observe that Python is the most common programming language required by users, followed by JavaScript. Such results might be motivated by the continuous adoption of these languages for Machine Learning and Web Development, respectively. The remaining programming languages present similar frequencies, even oscillating positions with each other, but with a large difference from the top two languages (placed at the bottom of the chart).
Analyzing the chart, we can observe a clear decline in posted questions after the release of ChatGPT for Python and JavaScript (top 2), the same observation made for general posted content analyzed in RQ3. However, for the remaining programming languages, there is no clear drop as they present a similar frequency of posted questions before and after the release. Statistically assessing whether there is a significant difference between the number of posted questions through the Wilcoxon rank sum test, we can reject the null hypothesis for all the evaluated programming languages (p-value < 0.01).
(a) Python (Programming Language)
(b) ReactJS (Framework)
(c) Pandas (Library)
Figure 8: Trend comparison of the number of questions associated with specific tags on Stack Overflow over time. Here, we just present the results for three tags; in our online appendix, we present the trend analysis for all evaluated tags.
Regarding the effect size, following the previous results reported in this RQ, we observe a moderate effect for most evaluated programming languages (seven out of 10) after the release of ChatGPT. However, when checking the effect size after the one-year milestone, we observe a large effect for all evaluated programming languages, revealing a substantial impact on question tagging these programming languages.
Different Impact on Most Cited Frameworks and Libraries
Figures 10 and  11 show the distribution of questions associated with the top ten most cited frameworks and libraries on Stack Overflow, respectively. Different from the chart for programming languages (Figure 9), we can observe different patterns regarding the distribution of questions here. Regarding the analysis of frameworks, we observe that ReactJS is the most cited framework, largely isolated from the others. Furthermore, that is the only framework in which we can clearly see a drop in the distribution of questions. Next, we have NodeJS and Android placed in the second and third positions, respectively. Although they oscillate in their positions, Android is reported in the second position for most of the evaluated time. Finally, the remaining frameworks are placed at the bottom of the chart, with some oscillations over time. Different from the programming languages domain, all the frameworks alternate positions except for the first one, showing the ranking of frameworks is more dynamic, with frequent shifts in their relative positions over time. Overall, it is also important to highlight the massive presence of frameworks based on JavaScript.
A similar distribution is also observed for the analysis of libraries. As highlighted in Figure 11, pandas is the most cited library, placed at the top of the chart with a large distance from the others. As previously discussed for the programming language Python, the adoption of pandas might be related to their adoption in Machine Learning, as it is commonly used for different tasks, like analyzing, cleaning, exploring, and manipulating data. Next, we observe a group of four libraries placed in the middle of the chart, while the remaining ones are placed at the bottom, oscillating positions over time. Finally, similar to the observations made for the framework analysis, here we observe a consistent number of libraries supporting different tasks on Machine Learning topics, showing the activity of users on these topics.
Aiming to statistically evaluate the impact on the number of posted questions, we also ran the Wilcoxon rank sum test. As a result, we observe different results when compared to our analysis of programming languages. Regarding the frameworks, we reject the null hypothesis for nine frameworks (p-value <0.05), except for Spring Boot (p-value = 0.06). A similar behavior was observed for the libraries under analysis, as we reject the null hypothesis for all of them (p-value <0.05), except ggplot2 (p-value = 0.34). Although a small fraction of subjects did not present a statistical drop in posted questions, when we analyze the posted content after the one-year milestone, we reject the null hypothesis for all evaluated frameworks and libraries (p-value <0.01). Regarding the effect size, similar to the previous results for programming languages, we observe a moderate effect for most evaluated libraries and frameworks (six out of 10) after the release of ChatGPT. However, when checking the effect size after the one-year milestone, we observe a large effect for all evaluated libraries and frameworks.
On the other hand, we also aim to have an initial insight into domains that exhibited increasing activity over time. For that, we first filtered the tags that appeared in all evaluated time windows of this study (before, after, and one year after ChatGPT ’s release). As a result, we identified 24 5642456424\,56424 564 commonly shared tags. Next, we examined the number of mentions associated with each tag, selecting those with a consistent increase over time, which resulted in 656 tags (2.68%). These tags span various domains, including new versions of programming languages and frameworks, tools, cloud services, and AI/ML-related technologies. Among these increasing tags, only four exhibited substantial overall activity, each surpassing 1000 mentions: openai-api, vercel, powerbi-desktop, and next-auth. These technologies have gained significant attention in recent years, with OpenAI API benefiting from the rise of LLMs, and Vercel and NextAuth growing in popularity alongside Next.js. However, it is important to highlight that the majority of the tags decreased over time, supporting our findings reported here.
RQ3.1 answer: Despite the observed general drop in posted content in Stack Overflow, initially, after the introduction of ChatGPT, we observe different impacts on the domains analyzed (programming languages, frameworks and libraries). Even analyzing the most cited topics from the previous domains, we observe that, in some cases, there is no statistical decline regarding the volume of posted content. However, after the one-year milestone following ChatGPT’s release, we observe a statistically significant drop in posted content regarding all domains analyzed.
4.4 How has the release of ChatGPT impacted the activity patterns of Stack Overflow users?
In this research question, we further investigate the impact of the release of ChatGPT on users’ activity on Stack Overflow. First, we discuss the overall impact on the user’s activity, and then, we delve into the challenges faced by users when recurring to LLMs to solve their questions.
@@ 4.4 How has the release of ChatGPT impacted the activity patterns of Stack Overflow users?
Active Users Pre-Release 314,179 62,007 67,240 239,045
@@
One-Year
Milestone
174,097 31,901 30,980 116,830
Inactive Users Post-Release 259,981 49,457 56,671 179,300
One-year
Milestone
259,207 48,112 50,823 171,771
Excl. Active Users Post-Release 239,613 43,967 46,876 146,785
One-year
Milestone
149,503 25,284 26,489 88,257
Rem. Users Post-Release 54,198 12,550 10,569 59,745
One-year
Milestone
24,594 6,617 4,491 28,573
Table 3: User activeness on Stack Overflow. The three first Active Users rows are related to users with active history on pre-, post, and one-year milestone of the ChatGPT’s release, respectively. The rows for Inactive, Exclusive Active, and Reminiscent Users are calculated by analyzing a specific window along with its preceding windows. For instance, post-release Inactive Users are identified as those who were reported as active during the pre-release window.
Overall Impact on User’s Activity Patterns in Stack Overflow
Table 3 presents the general observed frequency of users on Stack Overflow. Similar to our previous findings, we observe a general drop in active users (2nd and 3rd rows, respectively, in Table 3). Regardless of the type of user (questioner, and/or respondent, and commentator), all rates present a consistent drop. Such observation corroborates the drop in content posted on Stack Overflow (see Section 4.3). When we compare the number of users active after the release of ChatGPT, based on the different roles analyzed here, we observe that most users were not active before ChatGPT’s release(4th row in Table 3). Some may argue that these new active users could be associated with new accounts created after the release of ChatGPT. However, when checking the creation date of these accounts, most of them were created before ChatGPT release (see Table 4). Since different factors might influence these new active users, like sporadic users, it is important to highlight the commitment users have to their community. Such motivation might occur due to new technologies that are released and, consequently, further discussions about how to use and advocate for them.
Although we previously reported a drop in the number of posted questions on Stack Overflow, these active users posting their new questions contribute to new data being generated and, consequently, new versions of LLMs being trained on them. However, keeping users motivated and recruiting new ones to continue the knowledge propagation is a challenge that QA forums must deal with from now on. In Section 5, we further discuss the impact and the need for generating data and LLMs.
New Challenges faced by Users in Stack Overflow
Once we observe that current active users are new and reminiscent ones, we aim to investigate how these users deal with Stack Overflow combining with LLMs. For that, we investigate whether users replicate questions on Stack Overflow, which were previously asked on ChatGPT, checking whether the users explicitly mentioned that statement in the question content. First, we had to filter the questions asking for clarification about using ChatGPT (13%), like its API (18% and 7% for reminiscent and new users, respectively ). Further exploring the remaining questions, though users mention they have consulted ChatGPT for clarifications, 60% of the questions do not present the answers generated by ChatGPT (49% and 70% for reminiscent and new users, respectively). The main reason pointed out by the users is that the answers provided do not fix their problems, as they have tried different proposed suggestions. For example, when a user asks about constructing an imbalanced dataset based on a pre-defined gini-coefficient, they report that the ChatGPT could not even generate a working function: [^23] https://stackoverflow.com/questions/76008808/
{quoting}
I have been working on this problem for a few days now, but my programming and math skills are leaking behind. Also, ChatGPT was not able to construct a working function. If someone has any example code or formula I might use, this would be very helpful. Help is really appreciated!!
For the users that consider the suggestions provided by ChatGPT somehow useful, they leveraged them as a starting point for further discussions. For example, in one question, a user asks for support when using the Overpass API. [^24] https://stackoverflow.com/questions/75558305/ Based on the initial solution provided, the respondent analyzed it and provided their suggestion, which, based on the respondent, was adequately addressing the initial issue, while the solution provided by ChatGPT was syntactically invalid:
{quoting}
I suggest the following as a starting point, which isn’t exactly what you’ve asked for but way better than what ChatGPT offers as a solution (which is, in fact, invalid Overpass QL syntax).
In another question, when a user asks about converting datetime.datetime objects in timezone-aware datetime ones, ChatGPT provides its answer using the pytz library. [^25] https://stackoverflow.com/questions/74879121/ The respondent answers the question by providing code and using another library (pandas). Furthermore, the respondent also analyzes the answer provided by ChatGPT and warns the questioner about the deprecation of the pytz library. These scenarios highlight how the discussions started on ChatGPT are further extended in Stack Overflow; while questioners use generated answers to give context to their issues, respondents go further by exploring these answers and providing further feedback.
@@
Regarding the topics that represent challenges for Stack Overflow users, we observe that reminiscent and new users face the same challenges when trying to get answers from ChatGPT, and later appeal to Stack Overflow. Similar to RQ2 (see Section 4.2), we observe that questions regarding Frameworks and Libraries represent the most recurrent challenge faced by reminiscent and new users (41% and 38%, respectively), followed by question targeting Programming Languages (22% and 25%, for reminiscent and new users, respectively). Supporting our claims that these topics currently challenge LLMs, users know these limitations as they experience and recognize them in practice. For example, when a user asks for support regarding starting a new project using multiple frameworks like Next.JS, they consider the provided support of ChatGPT was not very helpful. They further argue that the failed ChatGPT’s attempt is due to their limited access to information dated until late 2021: [^26] https://stackoverflow.com/questions/75861694/
{quoting}
To be honest, the first thing I tried when I ran into walls was ChatGPT 4. It helped a little, but as it says in their documentation, it only knows up to late 2021. All the current libraries have changed […].
@@
RQ4 answer: After the introduction of ChatGPT, we observe a drop in the number of active users in Stack Overflow, even under different roles (questioners, respondents, and commentators). Most active users’ post-release were inactive before ChatGPT, showing that they are reminiscent users, as their accounts were created before the release. Finally, reminiscent and new users share the same challenges when addressing their questions on ChatGPT and later appealing to Stack Overflow.